# SwapMoE: Efficient Memory-Constrained Serving of Large Sparse MoE Models via Dynamic Expert Pruning and Swapping

### v2

@@ 3.4. Fine-grained Expert Profiling
(3) num_experts ={#$=\{\#experts_{1},\#experts_{2},...,\#experts_{L}\},$
@@
(4) s.t. $E_{\text{memory}}(\text{config})\leq{LIMIT}_{\text{memory}},$
@@
3.4.1. Memory Footprint and Latency Estimation
3.4.2. Accuracy Influence Modeling
Directly measuring the accuracy of a configuration on the target device is time-consuming because it requires running the MoE model under the configuration multiple times. When the configuration space is large, measurement is impractical. Predicting the accuracy is a promising option, but the complex correlation between the configuration and the accuracy makes it challenging, which have been studied in previous work (Zhang et al., 2021). While it is more convenient to predict accuracy indirectly through modeling, the accuracy of the model itself is inherently unexplainable.
To obtain $E_{\text{accuracy}}$, we first need to collect a small amount of profiling samples from the target device that can reflect the data distribution of the deployment scenario. Profiling samples must be labeled to measure the accuracy of SwapMoE under different configurations. Since the deployed models will be used in the target environment for a long time, it is feasible to collect such profiling data. The data labeling can be done manually or with an oracle model.
Next, we generate a set of random configurations of Virtual Experts. For each configuration, we use SwapMoE under the configuration to perform model inference with the profiling dataset. We collect the corresponding $E_{\text{accuracy}}$. Note that when collecting $E_{\text{accuracy}}$, SwapMoE is operated by the expert scheduler, which can refer to Section 3.2 and Section 3.3.
Finally, we use deep learning to fit the relation between Virtual Experts configurations and the model accuracy. We use a small DNN (two-layer linear network work) to model the relation because it is lightweight and sufficient. The model minimizes the residual sum of squares between the actual model $E_{\text{accuracy}}$ by the lightweight DNN. The training of this DNN is exceptionally fast, with very low computational requirements, typically completing in just a few minutes (on a Jetson Nano).
3.5. Profiling-guided Memory Planning
@@ 3.5.1. Expert I/O Frequency
3.5.2. Layer Space Allocation
To allocate the limited memory budget to different layers, allow layers with more memory to utilize more Virtual Experts, find the optimal configuration ($\hat{\text{config}}$) that maximizes the model accuracy while satisfying given memory constraints, we utilize expert planner to obtain the optimal memory allocation scheme based on the previously obtained performance model obtained from $E_{\text{accuracy}}$, $E_{\text{memory}}$, and $E_{\text{latency}}$ in Section 3.4.
@@ 3.5.2. Layer Space Allocation
Consequently, we decided to employ a genetic algorithm for the search process. Specifically, we initialize a set of configurations randomly, and iteratively update them based on their performance metrics estimated with $E_{\text{accuracy}}$, $E_{\text{memory}}$ and $E_{\text{latency}}$. In each iteration, we randomly change one parameter in each configuration and create new configurations by exchanging or averaging existing configurations. The configurations that violate resource constraints or yield suboptimal performance are removed. We observed that certain layers in the model may have a more significant impact on model performance. The genetic algorithm can perceive this characteristic and preserve experts that have a greater influence on performance. In Section 5, we will illustrate some configurations found by the algorithm.
4. Implementation
We implemented SwapMoE atop two popular MoE frameworks, Tutel (Hwang et al., 2022) for Swin-MoE, and Huggingface Transformers (Wolf et al., 2019) for Switch Transformer (Fedus et al., 2022) and GPTSAN (gpt, 2023), with 4k LoC added. Tutel uses customized operators for MoE layers, which employ token padding techniques to enhance load balancing and shape alignment between experts. Since the main focus of this paper is not to address the impact of padding, we directly adopted the token padding techniques used in Tutel.
Efficient implementation of Virtual Experts. To minimize the additional memory footprint caused by experts’ swapping in and out of main memory, we abstract Virtual Experts as a pointer concept. Virtual Experts remains stored in main memory, a portion of memory that is pre-allocated and does not require reallocation. The expert weights contained in Virtual Experts are those to which it points. When loading experts to Virtual Experts are necessary, we utilize the torch.copy_ API in PyTorch for transferring expert weights, thereby avoiding the need for additional main memory allocation. This process effectively updates the experts to which Virtual Experts points.
Genetic algorithm configurations. We provide the parameters used in the genetic algorithm to search the optimal Virtual Experts configuration. The population size is 50, the mutation rate is 0.5, the crossover rate is 0.01, and the number of generations is 5000. To ensure feasible results, we set the lower bound of the search space to 1 and the upper bound to the total number of experts in the MoE layer.
Profiling data processing. During the implementation of SwapMoE, we utilized 10% of the test data as profiling data. To have a sufficiently long profiling data set, we replicated this 10% of data multiple times and concatenated the beginning with the end. It is worth mentioning that label is not required for all the data in the online phase.
5. Evaluation
Figure 9. End-to-end latency achieved by SwapMoE and baselines under different memory budgets on Jetson AGX. SwapMoE achieves similar latency with the pruned compact models while with much more parameters in use.
@@ 5. Evaluation
Figure 10. The accuracy achieved by SwapMoE and the pruning baseline under different memory budgets on (a) Jetson AGX and (b) Jetson Nano. SwapMoE achieves significantly higher accuracy than the baseline in almost all cases.
We evaluate SwapMoE to answer the following questions:
  1. (1)
    Is SwapMoE able to provide better latency-accuracy and memory-accuracy tradeoffs than baselines?
  2. (2)
    Is SwapMoE able to find the optimal configuration that meets the constraints while maximizing accuracy?
  3. (3)
    What about the robustness of SwapMoE against different numbers of experts, input data locality, and hardware?
  4. (4)
    How much does each component contribute to the effectiveness of SwapMoE?
  5. (5)
    What about the overhead of SwapMoE?
5.1. Experimental Setup
Edge environments. We use two edge devices: a Jetson Nano and a Jetson AGX ORIN. The batch sizes are all set to 1 as in common edge-side continuous serving scenarios. We use different latency and memory budgets to simulate different resource constraints.
Tasks, datasets and models. We evaluate the performance of SwapMoE on three common DL tasks:
Object detection aims to detect objects in an image, predicting the object bounding boxes and categories. We chose Cascade with Swin-MoE (Liu et al., 2022) base backbone as the detection model, and trained 32 experts Swin-MoE models with the Shift dataset (Sun et al., 2022) for evaluation. The performance of the detection model is measured by mean average precision over Intersection over Union (mAP) with IoU threshold at 0.5, where higher mAP means better.
Summarization aims to summarize long texts into short texts. We select the most popular summarization model, Switch Transformer (Fedus et al., 2022) (16, 32, and 64 experts per layer, denoted as SwitchT-16, SwitchT-32 and SwitchT-64) and use the samsum dataset (Russakovsky et al., 2015) and report Rogue-2 accuracy, where higher Rogue-2 accuracy means better.
Language modeling aims to predict the next word given the previous words. We use GPTSAN (Artetxe et al., 2022) as the base model and use the Wikipedia-japanese (wik, 2023) dataset for evaluation. The performance of the language model is measured by perplexity, where lower perplexity means better.
The pre-trained weights of the Switch Transformers and GPTSAN are obtained from the official Huggingface Transformer repository. The Swin-MoE model were trained by ourselves according to the source code and the default configurations of their papers. All datasets can be downloaded from public websites.
Baselines. The basic baseline is original MoE model. To show the superiority of SwapMoE, we also compare it with the following baselines: ‘Pruning’ (keeping a certain portion of experts with the largest $||E_{i}||$ in each layer, no switching in/out), ‘On demand’ (Huang et al., 2023) (keep a certain portion of experts in main memory, load the requested expert from memory on demand).
5.2. Overall Runtime Performance
Method Constraints Achieved performance
Memory budget (GiB) Memory (GiB) Latency (s) mAP (%)
Original MoE \ 1.99 0.25 52.60
vExperts 0.9 0.94 0.13 51.23
1.2 1.11 0.15 51.65
1.5 1.44 0.19 52.26
1.8 1.79 0.23 52.58
(a) Swin-MoE
@@ 5.2. Overall Runtime Performance
(b) SwitchT-16
Table 1. Performance of SwapMoE with different MoE models under different memory constraints.
Furthermore, SwapMoE is a general method applicable to various tasks, including object detection, summarization, and language modeling tasks. In detection tasks, significant reductions in memory and latency resulted in minimal accuracy loss. However, in language modeling, reducing memory and latency introduced some deterioration in perplexity. This is because in computer vision (CV) tasks, the pixel values between consecutive input images are also continuous, thus exhibiting stronger locality. In contrast, for natural language processing (NLP) tasks, consecutive inputs are only semantically similar, such as in tone or knowledge domain, and therefore do not exhibit the same level of locality as in CV tasks.
5.3. Offline Planning Performance
SwapMoE can find optimal configurations that satisfy given constraints while maximizing model performance, as shown in Table 1. For example, in the object detection task, when given a memory budget of 1.5 GiB, the configuration found by SwapMoE allows the model to achieve an actual 1.44 GiB peak GPU memory usage and 0.19 s inference latency, satisfying the constraints while only experiencing a 0.32% decrease in accuracy compared to the original MoE model.
Furthermore, it is observed that as the constraints become looser, SwapMoE can achieve higher accuracy, and conversely, tighter constraints result in lower accuracy. It found optimal configurations under various constraints, respectively, to maximize the utilization of resources. For example, in the summarization task, when the memory budget changed from 3.5 GiB to 4 GiB, the actual peak GPU memory and inference latency of the model during runtime changed from 3.58 GiB and 0.76 s to 3.99 GiB and 0.80 s, respectively. The accuracy also changed from 0.23 to 0.24. However, our method is not always able to satisfy the budget. For example, in the summarization task with SwitchT-16, when the memory budget is set to 3.5 GiB, the actual peak memory usage and inference latency achieved by the model are 3.58 GiB and 0.76 s, respectively. This situation arises because the inference performance of the model itself is difficult to predict accurately (Zhang et al., 2021), leading to inaccurate performance modeling.
(a) SwitchT-16
(b) Swin-MoE
Figure 11. Distribution of allocated memory across layers under different memory budgets in SwapMoE.
Additionally, we find that the distribution of Virtual Experts across different layers differs in different tasks. In the language modeling task, SwapMoE tends to maintain more experts in the middle layers, as shown in Figure 11 (a). This suggests that for language models, the intermediate MoE layer has a greater impact on model inference, as the intermediate layers have a more significant influence on logical processing. In contrast, in the object detection task, SwapMoE maintains more experts in the early layers, as shown in Figure 11 (b). This indicates that for visual models, the foremost MoE layer has a greater impact on model inference, as the preceding layers have a more substantial influence on image feature extraction. For example, when the constraints are looser in the object detection task, the total number of SwapMoE increases, and the number of SwapMoE in the early layers increases. This indicates that the early layers are more important for the object detection task.
5.4. Robustness Analysis
Figure 12. Robustness of SwapMoE under different usage scenarios.
Refer to caption (a) Different data locality
(b) Different hardware
Refer to caption Refer to caption
Figure 12. Robustness of SwapMoE under different usage scenarios.
Figure 13. Ablation study: performance of SwapMoE with different components replaced.
@@ 5.4. Robustness Analysis
Different data locality. SwapMoE can be used with input data with different patterns. As shown in Figure 11(a), when SwapMoE works with lower temporal locality data (faster video with 3x video speed and random shuffled video frames), it is still able to find good performance-resource tradeoffs. And as the temporal locality of the data decreases, the performance of SwapMoE deteriorates. We observed that SwapMoE performs best on normal videos, followed by faster videos with 3x video speed, and finally random shuffle video frames. This is because as the temporal locality of the data decreases, the importance of experts in the MoE model becomes more unstable, resulting in poorer performance of SwapMoE under the same resource usage conditions. In other words, to achieve the same model performance, SwapMoE requires more resource usage on data with lower temporal locality.
Different hardware. SwapMoE can be used on devices with different I/O bandwidths. As shown in Figure 11(b), SwapMoE can find good performance-resource tradeoffs on the Jetson AGX device. Additionally, we implement two versions of SwapMoE on Jetson AGX. (1) CPU-GPU: swapping experts between CPU memory and GPU memory. (2) GPU-Storage: swapping experts between GPU memory and external storage. The experimental results demonstrate that despite lower storage bandwidth, SwapMoE is still able to find excellent resource-performance trade-offs. This is because SwapMoE only requires to swap a small number of experts for each sample at runtime. A small bandwidth between the main memory and the external memory is sufficient.
5.5. Ablation study
@@
(1) ‘Simple scheduling’: change the scheduling in SwapMoE component, remove amortized updating component, and calculate expert importance score where experts with more tokens to inference have higher scores. The perplexity of SwapMoE is better than ‘simple scheduling’ because the importance score obtained by expert scheduler is more accurate via approximating the outputs of experts, instead of simplely counting the number of tokens dispatched to each expert. Therefore, SwapMoE choose the important experts that are more suitable for the current input samples. Furthermore, ”simple scheduling” currently lacks amortized updating, which results in higher latency. This illustrates the necessity of amortized updating, as it can significantly reduce inference latency through the process of amortizing computations.
(2) ‘Simple planning’: change the planning in SwapMoE component to ‘simple planning’ and adopt an even Virtual Experts distribution, meaning that each MoE layer has an equal number of Virtual Experts, and then selects the best from these search spaces, where search space is very limited and does not encompass the optimal solution. The perplexity of SwapMoE is superior to ‘simple planning’ because the expert planner utilizes genetic search to find the optimal configuration for SwapMoE from a large configuration space. In contrast, ‘simple planning’ cannot find a worse configuration than SwapMoE due to the suboptimal search process.
@@
Table 2. The runtime overhead of SwapMoE.
5.6. Overhead
@@ 6. Related Work
Systems optimization for MoE model serving. Common techniques for optimizing MoE model serving include offloading and swapping memory. Huang et al. (Huang et al., 2023) propose to swap the experts from GPU memory to CPU memory to reduce the memory consumption of MoE models, incurring high latency overhead. SE-MoE (Shen et al., 2022) utilizes Ring Memory offloading to reduce GPU memory consumption of MoE models. However, these methods are not suitable for resource-constrained devices with dynamic latency and memory constraints. The proposed SwapMoE is the first work to optimize MoE model serving on resource-constrained devices with dynamic memory constraints.
@@
Optimization for dynamic DL model serving. MoE is a type of dynamic neural networks. Many approaches are introduced to enable or enhance such dynamic DL model serving in general. Nimble (Shen et al., 2021) is a system that optimizes, compiles, and executes dynamic neural networks on multiple platforms by using a dynamic type system and a lightweight virtual machine runtime. Model scaling approaches (Fang et al., 2018; Han et al., 2021; Wen et al., 2023) propose to adjust the model size/architecture on edge devices to meet different resource constraints. Remix (Jiang et al., 2021) proposes to use multiple models and dynamically switch between them during the inference process. Besides serving dynamic models, researchers have also attempted to slice the static model to dynamic components to achieve different resource-performance tradeoffs (Hou et al., 2022; Zhang et al., 2020). As compared with these approaches, our design is fundamentally different because it is based on the unique structure and characteristics of MoE.
@@
Efficient design of MoE models. Many existing approaches study the efficiency problem of MoE models from the model design perspective. GShard (Lepikhin et al., 2020) scales up Transformer with MoE and improves the quality and efficiency of multilingual machine translation. Task-MoE (Kudugunta et al., 2021) extract subnet from a large MoE model by a task-level routing strategy. Chen et al. (Chen et al., 2022) propose to prune non-professional experts according to the downstream task for efficient MoE deployment. MPoE (Gao et al., 2022) proposes to build a parameter-efficient MoE architecture by enforcing parameter sharing between the experts. AutoMoE (Jawahar et al., 2022) utilizes neural architecture search to automatically design more efficient MoE models. These methods need to modify and retrain the MoE model to reduce resource consumption. They are orthogonal to our approach and can use our method to further reduce resource consumption.
7. Conclusion
@@
This paper addresses the challenges of deploying MoE models to resource-constrained edge devices. We propose a framework called SwapMoE that adaptively reduces the inference costs of MoE models based on the memory constraints while preserving the model accuracy. Experimental results have demonstrated that our method can significantly reduce the inference costs of MoE models with reasonable accuracy degradation. Our approach creates a nice space of resource-accuracy trade-off of SoTA large MoE models.

### v3

@@ 3.3 Fine-grained Expert Profiling
num_experts ={#$=\{\#experts_{1},...,\#experts_{L}\},$ (2)
@@
s.t. $E_{\text{memory}}(\text{config})\leq{LIMIT}_{\text{memory}},$ (3)
@@
3.4 Profiling-guided Memory Planning
@@ 3.4.1 Expert I/O Frequency
3.4.2 Layer Space Allocation
To allocate the limited memory budget to different layers, allow layers with more memory to utilize more Virtual Experts, find the optimal configuration ($\hat{\text{config}}$) that maximizes the model accuracy while satisfying given memory constraints, we utilize memory planner to obtain the optimal memory allocation scheme based on the previously obtained performance model obtained from $E_{\text{accuracy}}$, $E_{\text{memory}}$, and $E_{\text{latency}}$ in Section 3.3.
@@ 3.4.2 Layer Space Allocation
Consequently, we employ the genetic algorithm Holland (1992) for the search process. Specifically, we initialize a set of configurations randomly, and iteratively update them based on their performance metrics estimated with $E_{\text{accuracy}}$, $E_{\text{memory}}$ and $E_{\text{latency}}$. In each iteration, we randomly change one parameter in each configuration and create new configurations by exchanging or averaging existing configurations. The configurations that violate resource constraints or yield suboptimal performance are removed. We observed that certain layers in the model may have a more significant impact on model performance. The genetic algorithm can perceive this characteristic and preserve experts that have a greater influence on performance. In Section 4, we will illustrate some configurations found by the algorithm.
4 Evaluation
4.1 Experimental Setup
Platforms. We use two devices: a Jetson Nano and a Jetson AGX ORIN. The batch sizes are all set to 1 as in common edge-side continuous serving scenarios. We use different latency and memory budgets to simulate different resource constraints.
Tasks, datasets and models. We evaluate the performance of SwapMoE on two common DL tasks:
Summarization aims to summarize long texts into short texts. We select the most popular summarization model, Switch Transformer Fedus et al. (2022) (16, 32, and 64 experts per layer, denoted as SwitchT-16, SwitchT-32 and SwitchT-64) and use the samsum dataset Russakovsky et al. (2015) and report Rogue-2 accuracy, where higher Rogue-2 accuracy means better.
Language modeling aims to predict the next word given the previous words. We use GPTSAN Artetxe et al. (2022) as the base model and use the Wikipedia-japanese wik (2023) dataset for evaluation. The performance of the language model is measured by perplexity, where lower perplexity means better.
The pre-trained weights of the Switch Transformers and GPTSAN are obtained from the official Huggingface Transformer repository. All datasets can be downloaded from public websites.
Baselines. The basic baseline is original MoE model. To show the superiority of SwapMoE, we also compare it with the following baselines: ‘Pruning’ (keeping a certain portion of experts with the largest magnitude $||E_{i}||$ in each layer, no switching in/out), ‘On demand’ Huang et al. (2023a) (keep a certain portion of experts in main memory, load the requested expert from memory on demand). ‘Pruning’ is akin to a simplified version of pruning-based methods without training, while ‘On demand’ is a simplified version of swapping-based approaches.
Figure 7: End-to-end latency achieved by SwapMoE and baselines under different memory budgets on Jetson AGX. SwapMoE achieves similar latency with the pruned compact models while with much more parameters in use.
@@ 4.1 Experimental Setup
Figure 9: The accuracy achieved by SwapMoE and the pruning baseline under different memory budgets on (a) Jetson AGX and (b) Jetson Nano. SwapMoE achieves significantly higher accuracy than the baseline in almost all cases.
4.2 Overall Runtime Performance
4.3 Offline Planning Performance
@@
Table 1: Performance of SwapMoE with SwitchT-16 under different memory constraints .
SwapMoE can find optimal configurations that satisfy given constraints while maximizing model performance, as shown in Table 1. For example, when given a memory budget of 2.0 GiB, the configuration found by SwapMoE allows the model to achieve an actual 1.64 GiB peak GPU memory usage and 0.82 s inference latency, satisfying the constraints while experiencing a 0.01 Rogue-2 score increase compared to the original MoE model.
Furthermore, it is observed that as the constraints become looser, SwapMoE can achieve higher accuracy, and conversely, tighter constraints result in lower accuracy. It found optimal configurations under various constraints, respectively, to maximize the utilization of resources. For example, in the summarization task, when the memory budget changed from 3.5 GiB to 4 GiB, the actual peak GPU memory and inference latency of the model during runtime changed from 3.58 GiB and 0.76 s to 3.99 GiB and 0.80 s, respectively. The accuracy also changed from 0.23 to 0.2. However, our method is not always able to satisfy the budget. For example, in the summarization task with SwitchT-16, when the memory budget is set to 3.5 GiB, the actual peak memory usage and inference latency achieved by the model are 3.58 GiB and 0.76 s, respectively. This situation arises because the inference performance of the model itself is difficult to predict accurately Zhang et al. (2021), leading to inaccurate performance modeling.
Figure 10: Distribution of allocated memory across layers under different memory budgets with SwitchT-16.
Additionally, we find that the distribution of Virtual Experts across different layers differs in different tasks. In the language modeling task, SwapMoE tends to maintain more experts in the middle layers, as shown in Figure 10 (a). This suggests that for language models, the intermediate MoE layer has a greater impact on model inference, as the intermediate layers have a more significant influence on logical processing.
4.4 Robustness Analysis
Figure 11: Ablation study: performance of SwapMoE with different components replaced.
@@ 4.4 Robustness Analysis
4.5 Ablation study
@@
(1) ‘Simple scheduling’: change the scheduling in SwapMoE component, remove amortized updating component, and calculate expert importance score where experts with more tokens to inference have higher scores. The perplexity of SwapMoE is better than ‘simple scheduling’ because the importance score obtained by our method is more accurate via approximating the outputs of experts, instead of simplely counting the number of tokens dispatched to each expert.
(2) ‘Simple planning’: change the planning in SwapMoE component to ‘simple planning’ and adopt an even Virtual Experts distribution, meaning that each MoE layer has an equal number of Virtual Experts, and then selects the best from these search spaces, where search space is very limited and does not encompass the optimal solution. The perplexity of SwapMoE is superior to ‘simple planning’ because the memory planner utilizes genetic search to find the optimal configuration for SwapMoE from a large configuration space. In contrast, ‘simple planning’ cannot find a worse configuration than SwapMoE due to the suboptimal search process.
@@
Table 2: The runtime overhead of SwapMoE.
4.6 Overhead
@@ 5 Related Work
Systems optimization for MoE model serving. Common techniques for optimizing MoE model serving include offloading and swapping memory. Huang et al. Huang et al. (2023a) propose to swap the experts from GPU memory to CPU memory to reduce the memory consumption of MoE models, incurring high latency overhead. SE-MoE Shen et al. (2022) utilizes Ring Memory offloading to reduce GPU memory consumption of MoE models. However, these methods are not suitable for resource-constrained devices with dynamic latency and memory constraints.
@@
Optimization for dynamic DL model serving. MoE is a type of dynamic neural networks. Many approaches are introduced to enable or enhance such dynamic DL model serving in general. Nimble Shen et al. (2021) is a system that optimizes, compiles, and executes dynamic neural networks on multiple platforms by using a dynamic type system and a lightweight virtual machine runtime. Model scaling approaches Fang et al. (2018); Han et al. (2021); Wen et al. (2023) propose to adjust the model size/architecture on edge devices to meet different resource constraints. Remix Jiang et al. (2021) proposes to use multiple models and dynamically switch between them during the inference process. Besides serving dynamic models, researchers have also attempted to slice the static model to dynamic components to achieve different resource-performance tradeoffs Hou et al. (2022); Zhang et al. (2020). As compared with these approaches, our design is fundamentally different because it is based on the unique structure and characteristics of MoE.
@@
Efficient design of MoE models. Many existing approaches study the efficiency problem of MoE models from the model design perspective. GShard Lepikhin et al. (2020) scales up Transformer with MoE and improves the quality and efficiency of multilingual machine translation. Task-MoE Kudugunta et al. (2021) extract subnet from a large MoE model by a task-level routing strategy. Chen et al. Chen et al. (2022) propose to prune non-professional experts according to the downstream task for efficient MoE deployment. MPoE Gao et al. (2022) proposes to build a parameter-efficient MoE architecture by enforcing parameter sharing between the experts. AutoMoE Jawahar et al. (2022) utilizes neural architecture search to automatically design more efficient MoE models. These methods need to modify and retrain the MoE model to reduce resource consumption. They are orthogonal to our approach and can use our method to further reduce resource consumption.
6 Conclusion
@@
This paper addresses the challenges of deploying MoE models to resource-constrained edge devices. We propose a framework called SwapMoE that adaptively reduces the inference costs of MoE models based on the memory constraints while preserving the model accuracy. Experimental results have demonstrated that our method can significantly reduce the inference costs of MoE models with reasonable accuracy degradation. SwapMoE creates a nice space of resource-accuracy trade-off of SoTA large MoE models. Our work does not have obvious ethical impacts, as we focusing on model inference acceleration.
7 Limitations
The tasks considered in this paper are relatively limited, and the proposed method has not been evaluated across a wide range of tasks. Subsequently, we will continue to expand our evaluation to other NLP tasks. The paper did not test larger MoE models (e.g., those exceeding 70B parameters) due to computational resource constraints. Instead, experiments were conducted only on smaller Switch Transformers models, demonstrating the effectiveness of the proposed method.
Acknowledgement
This work is supported by the National Natural Science Foundation of China (Grant No.62272261).
Appendix A Details of Fine-grained Expert Profiling
A.1 Memory Footprint and Latency Estimation
A.2 Accuracy Influence Modeling
Different distributions of Virtual Experts across layers may lead to different influences to model accuracy. To obtain the optimal configuration of expert distribution, we must have the ability to efficiently obtain the accuracy of each configuration. Directly measuring this accuracy influences on the target device is time-consuming because it requires running the MoE model under different configurations for multiple times. Since the configuration space of expert distributions is large, measurement is impractical. Therefore, we decide to use machine learning to model the accuracy influence based on profiling data.
Specifically, to obtain $E_{\text{accuracy}}$, we first need to collect a small amount of labeled profiling samples from the target device that can reflect the data distribution of the deployment scenario. The samples are then used to measure the accuracy of SwapMoE under different configurations. Since the deployed models will be used in the target environment for a long time, it is feasible to collect such profiling data. The data labeling can be done manually or with an oracle model.
Next, we generate a set of random configurations of Virtual Experts. For each configuration, we use SwapMoE under the configuration to perform model inference with the profiling dataset. We collect the corresponding $E_{\text{accuracy}}$. Note that when collecting $E_{\text{accuracy}}$, SwapMoE is operated by the runtime scheduler, which can refer to Section 3.1 and Section 3.2.
Finally, we learn the relation between Virtual Experts configurations and the model accuracy with a small DNN (containing two fully-connected layers with ReLU). The DNN is lightweight and sufficient. It minimizes the residual sum of squares between the actual accuracy and predicted $E_{\text{accuracy}}$. The training of this DNN is exceptionally fast, with very low computational cost, typically converged in just a few minutes (on Jetson Nano) with the prediction error less than 1%.